\begin{thebibliography}{ELMNP1}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\bibitem[BCHM]{bchm} C. Birkar, P. Cascini, C. Hacon, and J. McKernan,
\textit{Existence of minimal models for varieties of log general type},  J. Amer. Math. Soc. 23, no. 2, 405--468 (2010).

\bibitem[BH]{BH} C. Birkar and Z. Hu. \textit{Log canonical pairs with good augmented base loci}, Compos. Math. \textbf{150}, 579--592 (2014).

\bibitem[CD]{CD} S. Cacciola and L. Di Biagio, \textit{Asymptotic base loci on singular varieties}, Math. Z. \textbf{275} no. 1--2, 151--166 (2013).

\bibitem[CJK]{CJK} S. Choi, S. Jang, and D. Kim, \textit{Adjoint asymptotic multiplier ideal sheaves}, \href{https://arxiv.org/abs/2311.07441}{arXiv:2311.07441}, (2024).

\bibitem[CP]{CP} S. Choi and J. Park, \textit{Potentially non-klt locus and its applications}, Math. Ann. 366(1), 141--166 (2016).

\bibitem[ELMNP1]{ELMNP-base loci} L. Ein, R. Lazarsfeld, M. Musta\c{t}\u{a}, M. Nakamaye, and M. Popa, \textit{Asymptotic invariants of base loci}, Ann. Inst. Fourier (Grenoble) \textbf{56}(6) 1701--1734, (2006).

\bibitem[ELMNP2]{ELMNP-restricted volume} L. Ein, R. Lazarsfeld, M. Musta\c{t}\u{a}, M. Nakamaye, and M. Popa, \textit{Restricted volumes and base loci of linear series}, Amer. J. Math., 131(3): 607--651, (2009).

\bibitem[Fuj1]{Fuj-bpf} O. Fujino, \textit{Effective base point free theorem for log canonical pairs—Koll\'{a}r type theorem}, Tohoku Math. J. (2) \textbf{61} (2009), no. 4, 475--481.

\bibitem[Fuj2]{Fuj-book} O. Fujino, (2017). \textit{Foundations of the minimal model program} (Vol. 35). Tokyo: Mathematical Society of Japan.

\bibitem[Gon]{Gon} Y. Gongyo, \textit{On weak Fano varieties with log canonical singularities}, J. reine angew. Math. \textbf{665}, 237--252 (2012).

\bibitem[HH]{HH} K. Hashizume and Z. Hu, \textit{On minimal model theory for log abundant lc pairs}, J. Reine Angew. Math., \textbf{767} (2020), 109--159.

\bibitem[Les]{Les} J. Lesieutre, The diminished base locus is not always closed, Compos. Math. {\bf 150} (2014), no.~10, 1729--1741.

\bibitem[LT]{LT} V. Lazi\'{c} and N. Tsakanikas, \textit{On the existence of minimal models for log canonical pair}, Publ. Res. Inst. Math. Sci. \textbf{58}, (2022), no. 2, 311--339.

\bibitem[Nak]{Nak} N. Nakayama, (2004). \textit{Zariski-decomposition and abundance} (Vol. 14). Tokyo: Mathematical Society of Japan.

\bibitem[Nik]{Nik} V.V. Nikulin, \textit{A remark on algebraic surfaces with polyhedral Mori cone}, Nagoya Math. J. \textbf{157}, 73--92 (2000).

\bibitem[PS]{PS} Yu. G. Prokhorov and V.V. Shokurov, \textit{Towards the second main theorem on complements}, J. Algebraic Geom. 18 (2009), no. 1, 151--199.

\end{thebibliography}

\documentclass[a4paper,12pt]{amsart}

\makeindex

\usepackage{amsfonts,graphics,amsmath,amsthm,amsfonts,amscd,amssymb,amsmath,latexsym,euscript, enumerate}
\usepackage{epsfig}
\usepackage{flafter}
\usepackage[all,cmtip,line]{xy}
\usepackage{array}
\usepackage[english]{babel}
\usepackage{overpic}
\usepackage{subfig}
\usepackage{multirow}
\usepackage{microtype}
\usepackage{wrapfig}
\usepackage{ulem}

\usepackage[dvipsnames,svgnames,table]{xcolor}
\usepackage{graphicx}
\usepackage[hyphens]{url}
\usepackage{hyperref}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{proposition}[theorem]{Proposition}

\newtheorem*{theorem*}{Theorem}

\theoremstyle{plain}

\theoremstyle{definition} 
\newtheorem{definition}[theorem]{Definition}
\newtheorem{definition-lemma}[theorem]{Definition-Lemma}

\newtheorem{example}[theorem]{Example}

\newtheorem{remark}[theorem]{Remark}

\numberwithin{equation}{section}

\newcommand{\C}{\mathbb{C}}
\newcommand{\R}{\mathbb{R}}

\newcommand{\Q}{\mathbb{Q}}

\def\P{\mathbb{P}}

\DeclareMathOperator{\ord}{ord}

\newcommand{\bm}{\mathbf B_-}  
\newcommand{\B}{\mathbf B}
\newcommand{\bp}{\mathbf B_+}  
\def\pnklt{\operatorname{pNklt}}

\def\mult{\operatorname{mult}}

\def\Supp{\operatorname{Supp}}

\def\pnklt{\operatorname{pNklt}}
\def\nklt{\operatorname{Nklt}}
\def\nnef{\operatorname{NNef}}

\usepackage{mathtools}

\DeclarePairedDelimiterX{\inp}[2]{\langle}{\rangle}{#1, #2}
\DeclarePairedDelimiterX{\norm}[1]{\lVert}{\rVert}{#1}

\usepackage{mathtools}

\title[Anticanonical divisor with good asymptotic base loci]
{Anticanonical divisor with good asymptotic base loci}

\begin{document}

\author[S.~Choi]{Sung Rak Choi}
\author[S.~Jang]{Sungwook Jang}
\author[D.-W.~Lee]{Dae-Won Lee}
\address{Department of Mathematics, Yonsei University, 50 Yonsei-ro, Seodaemun-gu, Seoul 03722, Republic of Korea}
\email{sungrakc@yonsei.ac.kr}
\address{Center for Complex Geometry, Institute for Basic Science, 34126 Daejeon, Republic of Korea}
\email{swjang@ibs.re.kr}
\address{Department of Mathematics, Ewha Womans University, 52 Ewhayeodae-gil, Seodaemun-gu, Seoul 03760, Republic of Korea}
\email{daewonlee@ewha.ac.kr }

\thanks{S. Choi and D.-W. Lee are partially supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2302-03. D.-W. Lee is also partially supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (No. RS-2023-00237440 and 2021R1A6A1A10039823).}

\subjclass[2010]{Primary 14E30; Secondary 14J17, 14J45.}
\keywords{Anticanonical minimal model, Fano type variety, Potential pair}

\begin{abstract}
  In this paper, we give a characterization of Fano type varieties in terms of the asymptotic base loci of $-(K_X+\Delta)$. We also show that for a potentially lc pair $(X,\Delta)$, if no plc centers are contained in the augmented base locus $\bp(-(K_X+\Delta))$, then $(X,\Delta)$ has a good $-(K_X+\Delta)$-minimal model. This gives an analogous result of Birkar--Hu on the existence of good minimal models.
\end{abstract}

\maketitle



\section{Introduction}\label{sect:intro}

The minimal model program has been the most powerful tool in the classification of algebraic varieties. Several variations of the program have been developed, each serving its purpose in addressing the classification problem.
In the classical minimal model program, much attention has been paid to the study of birational modifications of varieties $X$ with respect to the canonical divisor $K_X$, especially in the case where $K_X$ is effective. It is well known that a remarkable breakthrough has been made by \cite{bchm} in the search for the minimal models of general type varieties. Varieties $X$ with effective anticanonical divisors $-K_X$ are also of importance. Among them, Fano varieties are the most well-studied classes of varieties due to their rich geometry with applications in various areas of mathematics. Fano type varieties are generalizations of Fano varieties and they share many similar properties with Fano varieties. By definition \cite[Lemma-Definition 2.6]{PS}, a Fano type variety $X$ is equipped with an effective divisor $\Delta$ such that $(X,\Delta)$ is a klt log Fano pair. It is known by \cite{bchm} that Fano type varieties are Mori dream spaces. Thus, we can run the $-K_X$-minimal model program on Fano type varieties and we expect to obtain Fano varieties as the final outcome. In particular, Fano type varieties can be considered as birational predecessors of Fano varieties. Thus, determining which varieties are of Fano type is an important task.

\medskip

We consider the following generalization of the notions of the singularities of pairs.
Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ pseudoeffective and $E$ a prime divisor over $X$. The potential discrepancy $\bar{a}(E;X,\Delta)$ at a prime divisor $E$ over $X$ is defined as $\bar{a}(E;X,\Delta)=a(E;X,\Delta)-\sigma_{E}(-(K_{X}+\Delta))$, where $a(E;X,\Delta)$ is the discrepancy of $(X,\Delta)$ at $E$ and $\sigma_{E}(-(K_{X}+\Delta))$ is the asymptotic divisorial valuation of $-(K_X+\Delta)$ along $E$. We say that the pair $(X,\Delta)$ is potentially klt (resp. potentially lc) if $\inf_{E}\bar{a}(E;X,\Delta)>-1$ (resp. $\ge -1$), where the infimum is taken over all prime divisors $E$ over $X$.  See Section \ref{sect:prelim} for the precise definitions.

It is proved in \cite{CJK} and \cite{CP} that a variety $X$ is of Fano type if and only if there exists an effective divisor $\Delta$ on $X$ such that $(X,\Delta)$ is potentially klt and $-(K_X+\Delta)$ is big. (See Section \ref{sect:prelim} for the definitions of potentially klt (pklt) and potentially lc (plc) pairs.) It is easy to see that in this characterization, we can allow $-(K_X+\Delta)$ to be pseudoeffective as long as $\Delta$ is big (Remark \ref{rem:pklt peff=FT}). We can also allow  $(X,\Delta)$ to be plc or lc as long as the plc centers or lc centers are in general position. The following is the main result of this paper.

\begin{theorem}\label{thrm:main}
Let $(X,\Delta)$ be a $\Q$-factorial plc pair with a klt variety $X$. Suppose that the following conditions are satisfied:
\begin{enumerate}[(1)]
    \item  $-(K_X+\Delta)$ is big,
    \item  no lc centers of $(X,\Delta)$ are contained in $\bm(\Delta)$, and
    \item  no plc centers of $(X,\Delta)$ are contained in $\bm(-(K_X+\Delta))$.
  \end{enumerate}
Then $X$ is of Fano type.
\end{theorem}

The three conditions (1), (2), (3) in Theorem \ref{thrm:main} are sharp in the sense that if any one of the conditions is removed, then the result fails. See Section \ref{sect:example} for the examples.


Birkar--Hu \cite[Corollary 1.2]{BH} proves that if $(X,\Delta)$ is an lc pair such that $K_X+\Delta$ is big and no lc centers of $(X,\Delta)$ are contained in $\bp(K_X+\Delta)$, then there exists a good minimal model of $(X,\Delta)$, thereby generalizing the main result of \cite{bchm}. This roughly shows that under such given conditions, the lc pair $(X,\Delta)$ can be treated as a klt pair since the minimal model program on $(X,\Delta)$ modifies the locus inside $\bp(K_X+\Delta)$. Inspired by this result, we also prove the following.

\begin{theorem}\label{thrm:main2-good anti min}
  Let $(X,\Delta)$ be a plc pair such that $-(K_X+\Delta)$ is a $\Q$-Cartier big divisor. If no plc centers of $(X,\Delta)$ are contained in $\bp(-(K_X+\Delta))$, then there exists a good $-(K_X+\Delta)$-minimal model.
\end{theorem}

Note that the varieties satisfying the conditions in Theorem \ref{thrm:main2-good anti min} are not necessarily of Fano type. We also note that if $\bp(-(K_X+\Delta))$ contains plc centers of $(X,\Delta)$, then there may exist a $-(K_X+\Delta)$-minimal model that is not good. See Example \ref{example:not 2}.

We expect that Theorem \ref{thrm:main2-good anti min} provides us with some hint toward $-K_X$-minimal model program since the $-K_X$-minimal model program works on the varieties satisfying the conditions in Theorem \ref{thrm:main2-good anti min} even though they are not Mori dream spaces. Exploring how such program works or establishing the flowchart for the program will require further deep investigation.

The rest of this paper is organized as follows: in Section \ref{sect:prelim}, we first recall the definitions of potential pairs and asymptotic base loci, and gather various ingredients for the proof of our main results. In Section \ref{sect:proofs}, we give the proofs of Theorems \ref{thrm:main} and \ref{thrm:main2-good anti min}. In Section \ref{sect:example}, we show by examples that the conditions of the main results are sharp.

\section{Preliminaries}\label{sect:prelim}
We work over the field $\C$ of complex numbers.

\subsection{Pairs}
A \textit{pair} $(X,\Delta)$ consists of a normal projective variety $X$ and an effective divisor $\Delta$ such that $K_{X}+\Delta$ is $\R$-Cartier. Let $f\colon Y\to X$ be a proper birational morphism with $Y$ normal projective variety. For some divisor $\Delta_Y$ on $Y$, we can write $K_{Y}+\Delta_{Y}=f^{\ast}(K_{X}+\Delta)$, and let $E$ be a prime divisor on $Y$. We say that such $E$ is a \textit{divisor over} $X$ and define the \textit{discrepancy} $a(E;X,\Delta)$ of $(X,\Delta)$ at $E$ as
$$ a(E;X,\Delta)=-\mult_{E}\Delta_{Y}, $$
where $\mult_{E}\Delta_{Y}$ denotes the coefficient of $E$ in $\Delta_{Y}$. A pair $(X,\Delta)$ is said to be \textit{kawamata log terminal} (abbreviated as \textit{klt}) if $\inf\limits_{E}a(E;X,\Delta)>-1$, where the infimum is taken over all prime divisors $E$ over $X$. If we replace the above strict inequality with inequality, i.e., $\inf\limits_{E}a(E;X,\Delta)\ge -1$, then we say that $(X,\Delta)$ is \textit{log canonical} (abbreviated as \textit{lc}).

\subsection{Potential discrepancy}
Let $Y$ be a smooth projective variety and $D$ a big $\R$-Cartier divisor on $Y$. For a prime divisor $E$ on $Y$, we define the \textit{asymptotic divisorial valuation} of $D$ along $E$ as $\sigma_{E}(D)=\inf\{\mult_{E}D' \mid 0\leq D'\sim_{\R}D\}$. If $D$ is a pseudoeffective $\R$-Cartier divisor, then we define as $\sigma_{E}(D)=\lim\limits_{\varepsilon\to 0^{+}}\sigma_{E}(D+\varepsilon A)$, where $A$ is an ample divisor on $Y$. The limit is independent of the choice of $A$ and it coincides with the definition for big $\R$-Cartier divisors. Moreover, for a given pseudoeffective $\R$-Cartier divisor $D$, there are only finitely many prime divisors $E$ such that $\sigma_{E}(D)>0$. See \cite{Nak} for more details.

Let $X$ be a normal projective variety and $D$ a pseudoeffective $\R$-Cartier divisor on $X$. If $E$ is a prime divisor over $X$, there exists a resolution  $f\colon Y\to X$ where $E$ is a prime divisor on $Y$. We define $\sigma_{E}(D)$ as $\sigma_{E}(D)=\sigma_{E}(f^{\ast}D)$. One can see that $\sigma_{E}(D)$ only depends on the valuation $\ord_{E}$ defined by $E$.

Now, we introduce the notion of potential pairs which was first introduced in \cite{CP}.

\begin{definition}
Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ pseudoeffective. For a prime divisor $E$ over $X$, we define the \textit{potential discrepancy} $\bar{a}(E;X,\Delta)$ of $E$ with respect to $(X,\Delta)$ as
$$ \bar{a}(E;X,\Delta)=a(E;X,\Delta)-\sigma_{E}(-(K_{X}+\Delta)). $$
If $\bar{a}(E;X,\Delta)>-1$ for all prime divisors $E$ over $X$, then we say that the pair is \textit{weakly potentially kawamata log terminal}.
We say that the pair $(X,\Delta)$ is \textit{potentially kawamata log terminal} (resp. \textit{potentially log canonical}) if
$$ \inf_{E}\bar{a}(E;X,\Delta)> -1\;(\text{resp.} \geq -1), $$
where the infimum is taken over all prime divisors $E$ over $X$.
\end{definition}

If $-(K_X+\Delta)$ is big, then the notions of weakly pklt and pklt coincide.

\begin{theorem}[{cf. \cite[Theorem 1.3]{CJK}}] \label{thrm:wpklt=pklt}
  Let $(X,\Delta)$ be a pair such that $-(K_X+\Delta)$ is a big $\Q$-Cartier divisor. Then the pair $(X,\Delta)$ is weakly pklt if and only if it is pklt.
\end{theorem}
\begin{proof}
  This is immediate if we let $D=-(K_X+\Delta)$ in \cite[Theorem 1.3]{CJK}.
\end{proof}

Let $E$ be a prime divisor over $X$, that is, there is a proper birational morphism $f\colon Y\rightarrow X$ with $Y$ normal projective variety and $E$ is a prime divisor on $Y$. We denote by $C_{X}(E)$, the \textit{center} $f(E)$ of $E$ on $X$.

\begin{definition}\label{def:pnklt locus}
Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ pseudoeffective. We define the \textit{potentially non-klt locus} $\pnklt(X,\Delta)$ of $(X,\Delta)$ as
$$ \pnklt(X,\Delta)=\bigcup C_{X}(E),$$
where the union is taken over all prime divisors $E$ over $X$ such that $\bar{a}(E;X,\Delta)\le -1$.

The \textit{non-klt locus} $\nklt(X,\Delta)$ is defined as the union $\bigcup C_X(E)$ for all prime divisors $E$ over $X$ such that $a(E;X,\Delta)\le -1$.
\end{definition}

In general, we do not know whether $\pnklt(X,\Delta)$ is closed or not. However, if $-(K_{X}+\Delta)$ is big, then $\pnklt(X,\Delta)$ is closed and it coincides with the non-klt locus of some pair.

\begin{proposition}[{\cite[Theorem 4.4]{CJK}}]\label{prop:CP complement}
Let $(X,\Delta)$ be a pair such that $-(K_{X}+\Delta)$ is a big $\Q$-divisor. Then there is an effective divisor $D$ such that $D\sim_{\Q}-(K_{X}+\Delta)$ and
$$ \pnklt(X,\Delta)=\nklt(X,\Delta+D). $$
\end{proposition}

\bigskip

If $(X,\Delta+D)$ is klt (resp. lc), then the effective divisor $D$ in Proposition \ref{prop:CP complement} is called a \textit{klt complement} (resp. \textit{lc complement}) of $(X,\Delta)$.

The following characterization of Fano type varieties is proved using Proposition \ref{prop:CP complement}.

\begin{theorem}[{cf. \cite[Corollary 4.10]{CJK}, \cite[Theorem 5.1]{CP}}]\label{thrm:pklt+big=FT}
  Let $X$ be a $\Q$-factorial normal projective variety. Then the following are equivalent:
  \begin{enumerate}[(1)]
    \item $X$ is a Fano type variety.
    \item There exists an effective $\Q$-divisor $\Delta$ such that $(X,\Delta)$ is pklt and $-(K_X+\Delta)$ is big.
  \end{enumerate}
\end{theorem}

\subsection{Asymptotic base loci}
Let $X$ be a normal projective variety and $D$ an $\R$-Cartier divisor on $X$. The \textit{stable base locus} of $D$ is defined as
$$ \B(D)=\bigcap\{\Supp(\Delta) \mid 0\le \Delta\sim_{\R}D\}. $$
The \textit{augmented base locus} of $D$ is defined as
$$ \bp(D)=\bigcap_{A}\B(D-A),$$
where the intersection is taken over all ample $\R$-divisors $A$.
The \textit{restricted base locus} of $D$ is defined as
$$ \bm(D)=\bigcup_{A}\B(D+A), $$
where the union is taken over all ample $\R$-divisors $A$. The restricted base locus $\bm(D)$ is in general not Zariski closed (\cite{Les}), whereas augmented base loci $\bp(D)$ are always Zariski closed. Note that $\bp(D)=\B(D-A)$ for any sufficiently small ample divisor $A$.

The \textit{non-nef locus} of $D$ is defined as
$$ \nnef(D)=\bigcup C_{X}(E),$$
where the union is taken over all the prime divisors $E$ over $X$ such that $\sigma_{E}(D)>0$.

In general, we have
$$ \nnef(D)\subseteq \bm(D)\subseteq \B(D)\subseteq \bp(D). $$

We can check that $D$ is nef if and only if $\nnef(D)=\bm(D)=\emptyset$. If $X$ is a smooth projective variety and $D$ a pseudoeffective $\R$-Cartier divisor on $X$, then we have $\nnef(D)=\bm(D)$ (cf. \cite{Nak}, \cite{ELMNP-base loci}). One can ask whether this equality also holds on a singular variety or not. 

\begin{proposition}[{\cite[Theorem 1.2]{CD}}]\label{prop:klt base}
Let $(X,\Delta)$ be a klt pair and $D$ a pseudoeffective $\R$-Cartier divisor on $X$. Then we have
$$ \nnef(D)=\bm(D). $$
\end{proposition}

Let $(X,\Delta)$ be a pair with $-(K_{X}+\Delta)$ pseudoeffective and $E$ a prime divisor over $X$. The center $C_{X}(E)$ of a prime divisor $E$ over $X$ such that $\bar{a}(E;X,\Delta)=-1$ is called a \textit{plc center} of $(X,\Delta)$. If $a(E;X,\Delta)=-1$, then the center $C_{X}(E)$ of $E$ on $X$ is called an \textit{lc center} of $(X,\Delta)$.

Suppose that $\bar{a}(E;X,\Delta)\le -1$. If $\sigma_{E}(-(K_{X}+\Delta))=0$, then we have $a(E;X,\Delta)=\bar{a}(E;X,\Delta)\le -1$. Hence, we can show that 
$$
\nklt(X,\Delta)\subseteq \pnklt(X,\Delta)\subseteq \nklt(X,\Delta)\cup \nnef(-(K_{X}+\Delta)),
$$
where the inequalities are strict in general (\cite[Lemma 4.1]{CP}). In particular, if $X$ is a klt variety, then Proposition \ref{prop:klt base} implies 
\begin{equation}\tag{$\#$}\label{nklt}
\nklt(X,\Delta)\subseteq \pnklt(X,\Delta)\subseteq \nklt(X,\Delta)\cup \bm(-(K_{X}+\Delta)).
\end{equation}

Using these inclusions, we prove the following proposition which plays a crucial role in the proof of Theorem \ref{thrm:main2-good anti min}.

\begin{proposition}\label{prop:lc complement}
  Let $(X,\Delta)$ be a plc pair such that $-(K_X+\Delta)$ is a big $\Q$-divisor. If no plc centers of $(X,\Delta)$ are contained in $\bp(-(K_X+\Delta))$, then there exists an lc complement $D$ of $(X,\Delta)$ such that $(X,\Delta+(1+\epsilon)D)$ is also an lc pair and $\nklt(X,\Delta+D)=\nklt(X,\Delta+(1+\epsilon)D)$ for a sufficiently small rational number $\epsilon>0$.
\end{proposition}
\begin{proof}
Since no plc centers of $(X,\Delta)$ are contained in $\bp(-(K_X+\Delta))$, we have $\nklt(X,\Delta)=\pnklt(X,\Delta)$. By Proposition \ref{prop:CP complement}, there is an effective divisor $D$ such that $D\sim_{\Q}-(K_{X}+\Delta)$ and
$$ \pnklt(X,\Delta)=\nklt(X,\Delta+D). $$
Furthermore, we can assume that $\Supp(D)$ does not contain any lc center of $(X,\Delta)$.
By \cite{ELMNP-base loci} and \cite{bchm}, we have for some ample $\Q$-divisor $A$
$$
\bp(-(K_{X}+\Delta))=\B(-(K_{X}+\Delta)-A).
$$
Let $Z_{1},\dots, Z_{r}$ be the lc centers of $(X,\Delta)$. By assumption, $Z_{i}\not\subseteq\bp(-(K_{X}+\Delta))$ for each $i$.
Therefore, there exists an effective divisor $D'$ such that $D'\sim_{\Q}-(K_{X}+\Delta)-A$ and $Z_{i}\not\subseteq\Supp (D')$ for all $i$. Since $A$ is ample, we can also find the required effective divisor $D$.

Let $f\colon Y\to X$ be a log resolution of $(X,\Delta+D)$ and $E$ a prime divisor on $Y$. Suppose that $C_{X}(E)\subseteq \nklt(X,\Delta)$, i.e., $C_X(E)$ is an lc center of $(X,\Delta)$. Since $\nklt(X,\Delta)=\nklt(X,\Delta+D)$, we have $a(E;X,\Delta+(1+\epsilon)D)=a(E;X,\Delta)\ge-1$ for any $\epsilon>0$. If $C_{X}(E)\not\subseteq \nklt(X,\Delta)=\nklt(X,\Delta+D)$, then $a(E;X,\Delta+D)>-1$ and $a(E;X,\Delta+(1+\epsilon)D)>-1$ for a sufficiently small $\epsilon>0$. Hence, $(X,\Delta+(1+\epsilon)D)$ is lc for a sufficiently small $\epsilon>0$. In particular, we also have $\nklt(X,\Delta+D)=\nklt(X,\Delta+(1+\epsilon)D)$ for a sufficiently small $\epsilon>0$.
\end{proof}

Now we recall the following result on the existence of minimal models of lc pairs.

\begin{theorem}[{\cite[Corollary 1.2]{BH}}]\label{thrm:BH}
Let $(X,\Delta)$ be an lc pair such that $K_X+\Delta$ is a big $\Q$-divisor and assume that no lc centers of $(X,\Delta)$ are contained in $\bp(K_X+\Delta)$. Then there exists a good minimal model of $(X,\Delta)$.
\end{theorem}

\begin{remark}\label{rem:two antiminmodels}
For an lc pair $(X,\Delta)$, there are two definitions of minimal models that differ slightly from each other. For a birational map $\varphi\colon X\dashrightarrow Y$, let $\Delta_Y\coloneqq\varphi_*\Delta$. The pair $(Y,\Delta_Y)$ is called a \textit{minimal model of} $(X,\Delta)$ (in the classical MMP sense) if the following conditions hold:
\begin{enumerate}
\item $\varphi$ is a contraction (i.e., there are no $\varphi^{-1}$-exceptional divisors),
\item $K_Y+\Delta_Y$ is nef and
\item $a(E;X,\Delta)\leq a(E;Y,\Delta_Y)$ holds for any prime divisors $E$ over $X$, where the inequality is strict if $E$ is $\varphi$-exceptional.
\end{enumerate}
It is expected that this model is obtained by running the minimal model program on the given lc pair \cite[Section 4.8]{Fuj-book}.

Another definition by Birkar--Shokurov is also considered often. For a birational map $\varphi\colon X\dashrightarrow Y$, let $\Delta_Y^{log}\coloneqq\varphi_*\Delta+\sum E_i$, where $E_i$ are $\varphi^{-1}$-exceptional prime divisors. The pair $(Y,\Delta_Y^{log})$ is called a \textit{minimal model of} $(X,\Delta)$ if
\begin{enumerate}
\item $(Y,\Delta_Y^{log})$ is a $\Q$-factorial dlt pair,
\item \textcolor{red}{$K_Y+\Delta_Y^{log}$} is nef and
\item $a(E;X,\Delta)\leq a(E;Y,\Delta_Y^{log})$ holds for any prime divisors $E$ over $X$, where the inequality is strict if $E$ is $\varphi$-exceptional.
\end{enumerate}
It is expected that this minimal model is obtained by running the minimal model program on a dlt blowup of the given lc pair. The minimal model in Theorem \ref{thrm:BH} is the minimal model in this sense.

We can run both minimal model programs and the respective flips exist. However, the terminations of the respective flips are still open problems. See \cite{Fuj-book} for more comprehensive details on both minimal model programs on lc pairs.

Despite the difference in the definitions, by \cite[Lemma 2.9(i)]{LT}, the existence of minimal models in both senses is equivalent. More precisely, there exists a minimal model in the classical MMP sense if and only if there exists a minimal model in the sense of Birkar--Shokurov. See also \cite[Theorem 1.7]{HH}.
\end{remark}

\begin{definition}\label{def:anti min model}
For a plc pair $(X,\Delta)$ and a birational map $\varphi\colon X\dashrightarrow Y$, let $\Delta_Y\coloneqq\varphi_*\Delta$. Then $\varphi\colon X\dashrightarrow Y$ (or the pair $(Y,\Delta_Y$)) is called a $-(K_X+\Delta)$-\textit{minimal model} if
\begin{enumerate}
  \item $\varphi$ is a contraction (i.e., $\varphi^{-1}$ has no exceptional divisors),
  \item $-(K_Y+\Delta_Y)$ is nef and
  \item $a(E;X,\Delta)\geq a(E;Y,\Delta_Y)$ for any prime divisor $E$ over $X$, where the strict inequality holds for any $\varphi$-exceptional divisor $E$ on $X$.
\end{enumerate}
If furthermore, $-(K_Y+\Delta_Y)$ is semiample, then $(Y,\Delta_Y)$ is called a \textit{good $-(K_X+\Delta)$-minimal model}.
\end{definition}

\begin{remark}
  A birational map $\varphi$ in Definition \ref{def:anti min model} satisfying the condition (3) is called a \textit{$-(K_X+\Delta)$-nonpositive map}. We note that the potential discrepancy $\overline{a}(E;X,\Delta)$ is preserved under a $-(K_X+\Delta)$-nonpositive birational map by \cite[Proposition 3.11]{CP}. Hence, the resulting model $(Y,\Delta_Y)$ in the above definition is necessarily an lc pair.
\end{remark}

\section{Proofs}\label{sect:proofs}

In this section, we give the proofs of the main results of this paper, Theorem \ref{thrm:main} and Theorem \ref{thrm:main2-good anti min}.
\smallskip

\begin{proof}[Proof of Theorem \ref{thrm:main}]
  If $(X,\Delta)$ is pklt, then $X$ is of Fano type by Theorem \ref{thrm:pklt+big=FT}. Therefore, we assume that $(X,\Delta)$ is not pklt.

  By the inclusions (\ref{nklt}) in Section \ref{sect:prelim}, we have the following inclusions
  $$\nklt(X,\Delta)\subseteq \pnklt(X,\Delta)\subseteq \nklt(X,\Delta)\cup \bm(-(K_X+\Delta)).$$
  Since $\pnklt(X,\Delta)\nsubseteq \bm(-(K_X+\Delta))$, we have $\pnklt(X,\Delta)=\nklt(X,\Delta)$ and hence $\pnklt(X,\Delta)$ consists only of lc centers of $(X,\Delta)$. Consequently, if $\Delta=0$, then $(X,\Delta)$ is a klt pair and $\nklt(X,\Delta)=\emptyset$. Thus, $(X,\Delta)$ is a pklt pair and $X$ is a Fano type variety by Theorem \ref{thrm:pklt+big=FT}. Therefore we may assume below that $\Delta\neq 0$.

  For a sufficiently small $\epsilon>0$, $-(K_X+(1-\epsilon)\Delta)=-(K_X+\Delta)+\epsilon\Delta$ is big. Therefore, by Theorem \ref{thrm:pklt+big=FT}, it suffices to prove that the pair $(X,(1-\epsilon)\Delta)$ is pklt. We prove that $\bar{a}(E;X,(1-\epsilon)\Delta)>-1$ for any prime divisor $E$ over $X$. This will imply that $(X,(1-\epsilon)\Delta)$ is pklt by Theorem \ref{thrm:wpklt=pklt}.

  First of all, since $X$ is a klt variety, the pair $(X,(1-\epsilon)\Delta)$ is klt so that $a(E;X,(1-\epsilon)\Delta)>-1$ for every prime divisor $E$ over $X$. Moreover, every lc center of $(X,\Delta)$ is contained in $\Supp(\Delta)$. Thus, for a prime divisor $E$ over $X$, if $C_X(E)$ is not contained in $\bm(-(K_X+(1-\epsilon)\Delta))$, then $\bar{a}(E;X,(1-\epsilon)\Delta)=a(E;X,(1-\epsilon)\Delta)>-1$.
  If $C_X(E)$ is contained in $\bm(-(K_X+(1-\epsilon)\Delta))$, then we prove $\bar{a}(E;X,(1-\epsilon)\Delta)>-1$ as follows. If $C_X(E)\subseteq \bm(-(K_X+\Delta))$, then $C_X(E)$ is not a plc center of $(X,\Delta)$ by the given condition (3). Hence, $\bar{a}(E;X,\Delta)>-1$. Moreover, we have the following inequalities
  \begin{align*}
    \bar{a}(E;X,(1-\epsilon)\Delta)&=a(E;X,(1-\epsilon)\Delta)-\sigma_E(-(K_X+(1-\epsilon)\Delta))\\
    &\geq a(E;X,\Delta)+\epsilon\mult_E f^*\Delta-\sigma_E(-(K_X+\Delta))-\epsilon\sigma_E(\Delta)\\
    &=\bar{a}(E;X,\Delta)+\epsilon(\mult_E f^*\Delta-\sigma_E(\Delta))\\
    &\geq \bar{a}(E;X,\Delta)>-1.
  \end{align*}

  If $C_X(E)\nsubseteq \bm(-(K_X+\Delta))$, then $\sigma_E(-(K_X+\Delta))=0$ and we have the following inequalities
  \begin{align*}
    \bar{a}(E;X,(1-\epsilon)\Delta)&=a(E;X,(1-\epsilon)\Delta)-\sigma_E(-(K_X+(1-\epsilon)\Delta))\\
    &\geq a(E;X,\Delta)+\epsilon\mult_E f^*\Delta-\sigma_E(-(K_X+\Delta))-\epsilon\sigma_E(\Delta)\\
    &= a(E;X,\Delta)+\epsilon(\mult_E f^*\Delta-\sigma_E(\Delta)).
  \end{align*}
  If $E$ is not an lc place of $(X,\Delta)$, then $a(E;X,\Delta)>-1$ and hence, $\bar{a}(E;X,(1-\epsilon)\Delta)>-1$. If $E$ is an lc place of $(X,\Delta)$, then $a(E;X,\Delta)=-1$. Note that by the condition $(2)$, we have $\sigma_E(\Delta)=0$. Furthermore, since every lc center of $(X,\Delta)$ is contained in $\Supp(\Delta)$, we have $\mult_E f^*\Delta-\sigma_E(\Delta)=\mult_E f^*\Delta>0$.
  Therefore, the inequality $\bar{a}(E;X,(1-\epsilon)\Delta)>-1$ holds.

  Hence, we have $\bar{a}(E;X,(1-\epsilon)\Delta)>-1$ for every prime divisor $E$ over $X$. Thus, the pair $(X,(1-\epsilon)\Delta)$ is pklt and $-(K_X+(1-\epsilon)\Delta)$ is big, which implies that $X$ is of Fano type by Theorem \ref{thrm:pklt+big=FT}.
\end{proof}

\begin{remark}\label{rem:pklt peff=FT}
  We note that if $(X,\Delta)$ is a $\Q$-factorial pklt pair with $-(K_X+\Delta)$ pseudoeffective and $\Delta$ big, then one can see that the pair $(X,(1-\epsilon)\Delta)$ is pklt and $-(K_X+(1-\epsilon)\Delta)$ is big. Hence, $X$ is a variety of Fano type by Theorem \ref{thrm:pklt+big=FT}.
\end{remark}

\begin{proof}[Proof of Theorem \ref{thrm:main2-good anti min}]
  By Proposition \ref{prop:lc complement}, for a sufficiently small rational number $\epsilon>0$, there exists an effective divisor $D$ such that $K_X+\Delta+(1+\epsilon)D\sim_{\Q} \epsilon D$ is big and the lc centers of $(X,\Delta+(1+\epsilon)D)$ coincide with the plc centers of $(X,\Delta)$. Since we have
  $$\bp(K_X+\Delta+(1+\epsilon)D)=\bp(D)=\bp(-(K_X+\Delta)),$$
  no lc centers of $(X,\Delta+(1+\epsilon)D)$ are contained in $\bp(K_X+\Delta+(1+\epsilon)D)$. By \cite[Corollary 1.2]{BH}, there exists a good minimal model $\varphi\colon X\dashrightarrow Y$ of $(X,\Delta+(1+\epsilon)D)$ in the sense of  Birkar--Shokurov. By Remark \ref{rem:two antiminmodels}, there also exists a minimal model $f\colon (X,\Delta+(1+\epsilon)D)\dashrightarrow (Y,\Delta_Y+(1+\epsilon)D_Y)$ in the classical sense, where $\Delta_Y\coloneqq f_*\Delta$ and $D_Y\coloneqq f_*D$.

  Now, we claim that $f\colon (X,\Delta)\dashrightarrow(Y,\Delta_Y)$ is a good $-(K_X+\Delta)$-minimal model. We note that by \cite[Corollary 5.6]{ELMNP-restricted volume}, $K_Y+\Delta_Y+(1+\epsilon)D_Y$ is log big, i.e., $(K_Y+\Delta_Y+(1+\epsilon)D_Y)\vert_W$ is big for any lc center $W$ of $(Y,\Delta_Y+(1+\epsilon)D_Y)$. Therefore, $K_Y+\Delta_Y+(1+\epsilon)D_Y$ is semiample by \cite[Theorem 1.1]{Fuj-bpf}. Thus, $(Y,\Delta_Y)$ is lc and $-(K_Y+\Delta_Y)$ is  semiample since
  $$K_Y+\Delta+(1+\epsilon)D_Y\sim_\Q K_Y+\Delta_Y+D_Y+\epsilon D_Y\sim_\Q \epsilon(-(K_Y+\Delta_Y))$$
  is semiample. The $-(K_X+\Delta)$-nonpositivity condition (3) can be checked easily as well.
\end{proof}

Theorem \ref{thrm:main2-good anti min} gives us no information on the existence of the (good) $-(K_X+\Delta)$-minimal models if some plc centers of $(X,\Delta)$ are contained in $\bp(-(K_X+\Delta))$. As of now, we are unaware of whether there exists a $-(K_X+\Delta)$-minimal model although we expect that we can still run partial $-(K_X+\Delta)$-MMP.

\section{Examples}\label{sect:example}

In this section, we prove by examples that the conditions in Theorems \ref{thrm:main} and \ref{thrm:main2-good anti min} cannot be weakened. First of all, it is easy to see that if $X$ is not a klt variety, then $X$ cannot be a Fano type variety.

\begin{example}[Condition (1) of Theorem \ref{thrm:main} is necessary, {\cite[Example 4.8]{CP}}]\label{example:not 1}
Let $S$ be a smooth rational surface as in $(\ast\ast\ast)$ of \cite[p.84]{Nik}. Then, $-K_S$ is nef, $\kappa(-K_S)=0$ and $\pnklt(S,0)=\emptyset$. Hence, the condition $(1)$ of Theorem \ref{thrm:main} is the only condition that is not satisfied and $S$ cannot be of Fano type since $-K_S$ is not big.
\end{example}

\begin{example}[Condition (2) of Theorem \ref{thrm:main} is necessary, {cf. \cite[Basic construction 5.1]{Gon}}]\label{example:not 2}
  Let $\pi\colon S\rightarrow \P^2$ be the surface obtained by blowing-up at $9$ very general points in $\P^2$ and $S\subset \P^N$ a projectively normal embedding. Let $\tilde{S}$ be the cone over $S$, $f\colon X\rightarrow \tilde{S}$ the blow-up at the vertex and $E$ the exceptional divisor of $f$. Moreover, since $-(K_X+E)$ is nef and big and $(X,E)$ is an lc pair, it is also a plc pair and hence
  $$\nklt(X,E)=\pnklt(X,E)=E.$$
  Also, one can see that $E\nsubseteq \bm(-(K_X+E))$ since $-(K_X+E)$ is nef and big. Note that $\bm(\Delta)=E$ is the only lc center, hence the condition $(2)$ in Theorem \ref{thrm:main} is the only condition that is not satisfied in this example. We know that $X$ is not of Fano type since $-(K_X+E)$ is not semiample.

\smallskip

Note that although the plc center $E$ is contained in $\bp(-(K_X+\Delta))$ and $E$ is not contained in $\B(-(K_X+\Delta))$ (as we will show below), there still exists a $-(K_X+\Delta)$-minimal model. However, the goodness of the $-(K_X+\Delta)$-minimal models in Theorem \ref{thrm:main2-good anti min} is not guaranteed.

\medskip

\noindent\textbf{Claim.} $E\subseteq\bp(-(K_X+\Delta))$

\medskip

    Since $-(K_X+E)$ is nef and big, we have
    $$\bp(-(K_X+E))=\mathrm{Null}(-(K_X+E)).$$
    Note that $E\cong S$ and we have the following equalities
    \begin{align*}
      -(K_X+E)^2\cdot E=(-(K_X+E)\vert_E)^2=(-K_E)^2=\left(-\pi^*K_{\P^2}-\sum_{i=1}^{9}E_i\right)^2=0,
    \end{align*}
    where $E_i$ are exceptional divisors over each point $p_i$.
    Hence, $E\subseteq \mathrm{Null}(-(K_X+E))$.

\medskip

\noindent\textbf{Claim.} $E\not\subseteq\B(-(K_X+\Delta))$

\medskip

    By adjunction formula and the isomorphism $E\cong S$, we have
    \begin{align*}
      h^0(E,-(K_X+E)\vert_E)=h^0(E,-K_E)&\cong h^0(S,-\pi^*K_{\P^2}-E_1-\cdots-E_9)\\
      &\cong h^0(\P^2, \mathcal{I}_{\{p_1,\dots,p_9\}}(3))\neq 0.
    \end{align*}
    On the other hand, we have the following short exact sequence.
    \begin{align*}
      0\rightarrow \mathcal{O}_X(-E)\rightarrow \mathcal{O}_X\rightarrow \mathcal{O}_E\rightarrow 0
    \end{align*}
    By tensoring $\mathcal{O}_X(-(K_X+E))$, we obtain the following long exact sequence.
    \begin{align*}
      0&\rightarrow H^0(X,-(K_X+E)-E)\rightarrow H^0(X,-(K_X+E))\rightarrow H^0(E, -(K_X+E)\vert_E)\\
      &\rightarrow H^1(X,-(K_X+E)-E)\rightarrow\dots
    \end{align*}
    Since $-(K_X+E)-E=K_X-2(K_X+E)$ and $-(K_X+E)$ is nef and big, one can apply Kawamata--Viehweg vanishing theorem and obtain $H^1(X,-(K_X+E)-E)=0$. This gives us a surjection $H^0(X,-(K_X+E))\rightarrow H^0(E,-(K_X+E)\vert_E)\neq 0$. Since this restriction map is not a zero map, $E$ is not contained in the base locus of $\lvert -(K_X+E)\rvert$. Hence, we have $E\nsubseteq \B(-(K_X+E))$.

\end{example}

\begin{example}[Condition (3) of Theorem \ref{thrm:main} is necessary]\label{example:not 3}
  We use the same notations from Example \ref{example:not 2}. The pair $(X,(1-\epsilon)E)$ for sufficiently small $\epsilon$ satisfies all the conditions except $(2)$. In this case, we have $\nklt(X,(1-\epsilon)E)=\emptyset$, $\pnklt(X,(1-\epsilon)E)=\bm(-(K_X+(1-\epsilon)E))=E$.
\end{example}


\end{document}